OpenStudy (anonymous):

33. What are the minimum, first quartile, median, third quartile, and maximum of the data set? 18, 20, 11, 10, 8, 6, 12, 4

OpenStudy (kropot72):

First put the data in ascending order: 4, 6, 8, 10, 11, 12, 18, 20 Now you can easily find the minimum and maximum. The median is the value halfway between the two values either side of themiddle, \[\frac{1}{2}\times (10+11)=\] Next partition the data into two equally sized sub-sets: {4, 6, 8, 10} and {11, 12, 18, 20} The lower quartile is the median of the first of the two sub-sets, \[\frac{1}{2}\times (6+8)=?\] The upper quartile is the median of the second of the two sub-sets,\[\frac{1}{2}\times (12 +18)=\]

OpenStudy (anonymous):

THANK YOU :D

OpenStudy (kropot72):

You're welcome :)